Density of substance at 0°C is 10gm/cc and at 100°C it`s density is 9.7 gm/cc . The coefficient of linear expansion of the substance is

Grade:11

1 Answers

Vikas TU
6 years ago
Dear Student,
We kne that the formulae for
Coefficient of volume expansion:
Y = Δρρ.ΔT)
=(10−9.7)10×(100−0
)=3×10−4 .
Subsequently, coefficient of straight extension 
Y3=10−4/∘C.
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)